A certain mass of a gas containing 43.5cm³ at 15°C is heated at constant pressure to 85°C. It then occupies a volume of 54cm³. Calculate the cubic expansivity.


1
Expert's answer
2021-01-07T10:55:20-0500

The cubic expansivity shows how the change in volume relates to the initial volume per unit of temperature and can be expressed as


"\gamma=\frac{V_2-V_1}{V_1(\theta_2-\theta_1)},\\\space\\
\gamma=\frac{54-43.5}{43.5(85-15)}=3.45\space\frac{1}{\text{°C}}."